I need help trying to solve this problem. Thank you in advance, for helping me out. I really appreciate your time and efforts. Have a terrific day :) Here is the question? Factor each trinomial:
c^2- 9c + 8

Found 2 solutions by josgarithmetic, lpcruz:
Answer by josgarithmetic(39800)

Find two integers whose product is 8 and whose sum is NEGATIVE 9. Both integers will be negative (because their product is shown resulting in POSITIVE 8).

They are -1 and -8.

Answer by lpcruz(7)

c^2 9c+8
pq=8
p+q=-9
-1, -8
Final Answer= (c-1)(c-8)